OpenStudy (anonymous):

Two lines are parallel. The first line passes through (6, 1), and its y-intercept is (0, –11). Which of the following cannot be the equation of the second line?

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

What do you know about parallel lines?

OpenStudy (whpalmer4):

in particular, what is the relationship between their slopes?

OpenStudy (anonymous):

not much really so im trying to find someone that could tutor and help me on certain subjects

OpenStudy (anonymous):

well both of the lines are parallel which means they both are positive and have the same slope

OpenStudy (whpalmer4):

that's what I'm looking for, good! Okay, we have the first line going through (6,1) and (0,-11). Can you find the slope from that?

OpenStudy (anonymous):

yes hold on let me graph it

OpenStudy (whpalmer4):

By the way, you haven't posted the answer choices yet, so it will be a bit awkward to get you all the way through this problem, given that you are supposed to identify which of the following CANNOT be an equation of the line...

OpenStudy (anonymous):

ok hold on

OpenStudy (whpalmer4):

actually, you don't need to graph to get the slope... \[m = \frac{y_2-y_1}{x_2-x_1}\]where the line goes through points \((x_1,y_1)\) and \((x_2,y_2)\)

OpenStudy (anonymous):

A. B. y = 2x + 8 C. D. y = –2x – 8

OpenStudy (anonymous):

C.

OpenStudy (anonymous):

ok c wont show

OpenStudy (anonymous):

-12/-6

OpenStudy (anonymous):

thats the slope

OpenStudy (whpalmer4):

looks like A didn't show either. can you reduce that fraction at all?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

thats the reduced fraction

OpenStudy (whpalmer4):

yes. so the slope of our first line is 2. What will be the slope of the other line?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

that make the answer o this problem b

OpenStudy (whpalmer4):

no, we need to know which equation CANNOT be the other line. These lines are parallel, so they have the same slope. If you see an equation which has a different slope, THAT's the one that is the correct answer choice, because it cannot be parallel if it has a different slope.

OpenStudy (anonymous):

ok well d has a different slope with -2x

OpenStudy (anonymous):

and a is -1/2 and c is 1/2

OpenStudy (whpalmer4):

I'm sorry, can you type the equations of the answer choices?

OpenStudy (anonymous):

ok and srry i put the wrong answer choices

OpenStudy (anonymous):

A. y = 2x + 7 B. 2x + y = 6 C. 4x – 2y = 2 D. 6x = 3y – 2

OpenStudy (whpalmer4):

Ah, they make us work a little harder :-) Can you rearrange all of those so they are in the form y = mx + b? A doesn't need any changes, but the other 3 do.

OpenStudy (anonymous):

im not sure how to do that though

OpenStudy (whpalmer4):

Sure you do. Just solve each of those equations for y.

OpenStudy (whpalmer4):

I'll do B: \[2x+y = 6\]subtract \(2x\) from each side \[2x-2x+y=6-2x\]\[y = -2x+6\]

OpenStudy (whpalmer4):

C. \[4x-2y = 2\]\[-2y = 2-4x\]\[y = -1+2x = 2x - 1\]

OpenStudy (whpalmer4):

D. \[6x=3y-2\]\[6x+2=3y\]\[y=2x+\frac{2}{3}\]

OpenStudy (whpalmer4):

So, which of those 4 equations cannot be a parallel line to our first line?

OpenStudy (anonymous):

hold on i want to right down how you did it so i know how the next time

OpenStudy (anonymous):

and b

OpenStudy (whpalmer4):

Yes, b is the correct answer choice, because it is the only one with a different slope than m=2.

OpenStudy (anonymous):

yes thank you